OpenStudy (anonymous):

PLEASE HELP WILL AWARD FAN AND MEDAL!!!!!! circle a as an area of 254.469 square units. What is the circumference of circle a ? Round your answer to the nearest hundredth, if necessary. A. 9.00 units B. 15.95 units C. 56.55 units D. 81.00 units

OpenStudy (anonymous):



OpenStudy (anonymous):

and so that is the opposite of \[\pi \times R ^ 2\]

OpenStudy (anonymous):

yes and i would plug in 254.469 for area correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but not for circumference.

OpenStudy (anonymous):

well wouldnt you solve for the radius then plug it in to the formula to find circumference

OpenStudy (anonymous):

and \[\pi = 3.1415926535898\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes the formula for the circumference of a circle is R x \[\pi \times 2\]

OpenStudy (anonymous):

okay so then i solve for r then plug that in and find the circumference?

OpenStudy (anonymous):

yes, that is right

OpenStudy (anonymous):

plus i am only in grade 3

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can you medal and fan now?

OpenStudy (anonymous):

medal you and fan

OpenStudy (anonymous):

is c correct?

OpenStudy (anonymous):

thank you soo much :)